Gingerbread cookies become inedible if not eaten quickly enough. Clarence is trying to determine how many cookies to bake for each day. The probabilities are below

Probability

30

35

40

45

50

Value

.2

.25

.1

.2

.25

Cookie cost c=$2
Selling price p=$4
Salvage value s=$0.30

How many should he bake? Using a probability cost chart, calculate the likely profit for this venture as well.
Answer #1

Formula

Profit = selling price-cost price = 4-2 = 2

Cost of overstocking = cost-salvage value = 2-0.3 = 1.7

We can see baking quantity 40 yields highest expected profit 67.975 so 40 should be baked